OpenStudy (anonymous):

lim (4sinx-4)/(lnsinx) as x approaches pi/2?

OpenStudy (anonymous):

Are you familiar with derivative?

OpenStudy (anonymous):

i'm just quite confused in \[\frac{ \pi }{ 2 }\] because it's the same with 90 right?

hartnn (hartnn):

yup, pi/2 = 90

OpenStudy (anonymous):

i use a calculator to solve sin 90 and sin pi/2, how come the ans is different?

hartnn (hartnn):

because your calculator can be only one of the mode, either in degree mode or in radian mode.

hartnn (hartnn):

if u got sin 90 =1 then your calcy is in degree mode

OpenStudy (anonymous):

pi/2 RADIANS = 90 DEGREES

OpenStudy (anonymous):

What is your question. Is that pi/2 or the imit?

OpenStudy (anonymous):

can I have an opinion on what to do?

hartnn (hartnn):

are u aware of L'Hopital's Rule ?

OpenStudy (anonymous):

yep

hartnn (hartnn):

can u differentiate numerator and denominator ? and tell me what u get.

hartnn (hartnn):

*differentiate numerator and denominator separately

OpenStudy (anonymous):

\[\frac{ 4cosx }{ \frac{ cosx }{ sinx } }\]

hartnn (hartnn):

very good :) that would be 4 sin x , isn't it ?

hartnn (hartnn):

@lambchamps did u get how it is 4 sin x ??

OpenStudy (anonymous):

i got it, thanks for the reminder

hartnn (hartnn):

ok, so did u get final answer as 0 ?

OpenStudy (anonymous):

with \[\frac{ 4cosx }{ \frac{ cosx }{ sinx } }\] cosx would be cancelled right and the sinx would be a numerator so it is now 4sinx, here i got the answer 4

hartnn (hartnn):

ok, thats correct, i just wanted to confirm that u got 4, thats why i asked whether u got 0.